(d) A representative consumer has a utility function U (x, y) = xy. A representative firm makes

good x and has a production function x = f(k, l) = (kl)0.25and an unavoidable fixed cost

equal to A. There are 100 consumers and, initially, 100 firms. Prices are w = v = Py = 1 and Px is determined in a competitive market. Representative consumer income is I = 2.

(d) Solve for the equilibrium price and quantity of good x?
Answer #1

OCX, Y 2 = xy Du = mox ox = y ou = muy = x - Oy ų MRS = mux = muy BL: Pxx+ Pyy = I - At optimal choice mRs = put y = (Px) x ix = = fck, D x (ke) 0.25 x0.25 0.25 Ox = mpe = 0.25%0-25 20.25-1 o e mpe= 0.25k 0.25 e 0.75 pxmpx = 0.25 x 0.25-) 20-25 ok MPc= 0x2 + 12x2 c= 2x2 de = nec = 4x O suppey curve for an firm is given by individual p= mc Рx = Tмск 4x Px = x = Px since Thu